OpenStudy (chupacabraj):

A scalar potential U in one dimension is given by U = 1/2 kx^2, where k is a constant and x position. Find the conservative force magnitude F for this potential.

OpenStudy (chupacabraj):

I thought force is the derivative of U?

OpenStudy (agent0smith):

It is. That's the spring potential energy formula. The spring force is F = kx.

OpenStudy (chupacabraj):

That's what I picked on my exam but I got it wrong lol

OpenStudy (chupacabraj):

These were the answer choices: F=kx, F= -kx, F=kx^2, F= 1/2kx^2, and F= GM m/x^2

OpenStudy (chupacabraj):

I chose F= kx but apparently it's wrong? Why?

OpenStudy (agent0smith):

Technically it's F = -kx since it's a restoring force, butttt the question says "conservative force magnitude F" keyword being magnitude.

OpenStudy (chupacabraj):

hehe so I should've gotten the points for this question then